# Classification and stability of penalized pinned elasticae

### v1

@@ 1. Introduction
where $k$ and $s$ respectively denote the signed curvature and the arclength parameter of $\gamma$. A critical point of the bending energy under the length-constraint is called Euler’s elastica, and it is known as a model of an elastic rod (cf. [2, 9, 13]).
@@
where $\lambda>0$ is called a penalization parameter, and $L[\gamma]=\int_{\gamma}\mathrm{d}s$, which is an object of interest in the recent literature (see e.g. [1, 4, 10, 14, 16, 23, 24, 25] for studies of critical points, [5, 6, 7, 8, 15, 22] for gradient flows, and references therein).
@@
$W^{2,2}_{\rm imm}(0,1;\mathbf{R}^{2}):=\Set{\gamma\in W^{2,2}(0,1;\mathbf{R}^{2})}{\,|\gamma^{\prime}(x)|\neq 0\ \ \text{for all}\ x\in[0,1]}.$
@@
Inspired by this question, we will obtain various properties of penalized pinned elasticae, such as (i) complete classification, (ii) stability results, (iii) uniqueness of nontrivial minimizers, and (iv) consequences for the elastic flow.
@@
Suppose that $\gamma\in A_{\ell}$ is a critical point of $\mathcal{E}_{\lambda}$ in $A_{\ell}$. Then, $\gamma$ is a trivial line segment; otherwise, up to reflection and reparametrization, $\gamma$ is represented by one of the following:
@@
Explicit formulae for each penalized pinned elasticae are also obtained (see Definition 2.7 below). Note that the classification strongly depends on the relation of $\lambda\ell^{2}$ and $\hat{\lambda}$ defined in (1.4). More precisely, if $\lambda\ell^{2}<\hat{\lambda}$, then $n_{\lambda,\ell}=1$ holds so that $\gamma_{\rm sarc}^{\lambda,\ell,1}$ and $\gamma_{\rm larc}^{\lambda,\ell,1}$ appear. On the other hand, if $\lambda\ell^{2}=\hat{\lambda}$, then $\gamma_{\rm sarc}^{\lambda,\ell,1}$ and $\gamma_{\rm larc}^{\lambda,\ell,1}$ coincide, and if $\lambda\ell^{2}>\hat{\lambda}$, then $n_{\lambda,\ell}\geq 2$ holds so that $\gamma_{\rm sarc}^{\lambda,\ell,1}$ and $\gamma_{\rm larc}^{\lambda,\ell,1}$ do not appear (see Figure 2). Such ‘bifurcation phenomena’ for elasticae have previously been found in [12, Section 2.3.3-2.3.5].
@@
We also seek to compare the energies of the above critical points and study which one is energy-minimal (if one excludes the trivial line segment). As the explicit formulae show, $\mathcal{E}_{\lambda}[\gamma_{\rm larc}^{\lambda,\ell,1}]$ and $\mathcal{E}_{\lambda}[\gamma_{\rm loop}^{\lambda,\ell,1}]$ converge to 00 as $\lambda\to 0$. This implies that for small $\lambda>0$ it is not easy to determine which has less energy, $\gamma_{\rm larc}^{\lambda,\ell,1}$ or $\gamma_{\rm loop}^{\lambda,\ell,1}$. On the other hand, for larger $\lambda>0$ a comparison with $\gamma_{\rm sarc}^{\lambda,\ell,1}$ is also required. Indeed, since
@@
Here we are interested in the question of eventual embeddedness. We find a sharp energy threshold below which the above flow destroys any self-intersection in finite time. For more results on (not necessarily eventual) embeddedness of elastic flows we refer to [3, 17].
@@
The value $C_{\lambda,\ell}$ is optimal in the sense that there is a smooth curve $\gamma_{0}$ such that $\mathcal{E}_{\lambda}[\gamma_{0}]=C_{\lambda,\ell}$ and the $\lambda$-elastic flow with initial datum $\gamma_{0}$ is not embedded for all $t\geq 0$. Actually, one may choose $\gamma_{0}=\gamma_{\rm loop}^{\lambda,\ell,1}$. The instability of $\gamma_{\rm loop}^{\lambda,\ell,1}$ in Theorem 1.2 allows us to construct a nonembedded smooth curve $\gamma_{0}\in A_{\ell}$ satisfying the condition (1.9), cf. Remark 5.3. From Theorem 1.4 we then infer that the flow destroys each self-intersection of $\gamma_{0}$ in finite time.
We close this introduction by comparing the properties of our penalized pinned elasticae with those already known from [19] about the critical points of $B$ in
@@
where $L>\ell$ is given. It turns out that the curves of Theorem 1.1 are also critical points in $A_{\ell,L}$ if $L$ is chosen to be their corresponding length. In [19], stability of these curves in $A_{\ell,L}$ is investigated. A major difference is that depending on the choice of $L$ it is possible that $\gamma_{\rm sarc}^{\lambda,\ell,1}$ is stable in $A_{\ell,L}$, whereas this is never the case for our penalized problem (cf. Theorem 1.2). Another difference is that in $A_{\ell,L}$ stable critical points are always unique whereas for our penalized problem the number of stable critical points depends on $\lambda$. Exposing these differences to the fixed-length problem is a main novelty of the present article.
@@ Acknowledgments
This work was initiated during the second author’s visit at Freiburg University. The second author is very thankful to the first author for his warm hospitality and providing a fantastic research atmosphere. Both authors are grateful to Tatsuya Miura for helpful suggestions. The second author is supported by FMfI Excellent Poster Award 2022 and JSPS KAKENHI Grant Number 24K16951.
@@ 2. Classification
    The first derivative vanishes for any perturbation of the kind $\gamma+\varepsilon\eta$ with $\eta\in W^{2,2}(0,1;\mathbf{R}^{2})\cap W^{1,2}_{0}(0,1;\mathbf{R}^{2})$ one has
    (2.1) $\frac{d}{d\varepsilon}\mathcal{E}_{\lambda}[\gamma+\varepsilon\eta]\Big{|}_{\varepsilon=0}=0.$
@@
First we show that $k$ is smooth on $(0,L)$. For arbitrary $\varphi\in W^{2,2}(0,L;\mathbf{R}^{2})\cap W^{1,2}_{0}(0,L;\mathbf{R}^{2})$ and consider $\eta(x):=\varphi(s(x))$ for $x\in[0,1]$, where $s$ denotes the arclength function, i.e., $s(x):=\int_{0}^{x}|\gamma^{\prime}|$. Inserting $\eta$ into (2.1), and using the known formulae of first derivative of $B$ and $L$ (cf. [18, Lemma A.1]), we deduce that
@@
where $\mathbf{t}$ and $\mathbf{n}$ denote the unit tangent vector and the unit normal vector of $\tilde{\gamma}$, respectively. Since this identity holds in particular for all $\varphi\in C^{\infty}_{\rm c}(0,L;\mathbf{R}^{2})$, we can deduce from a standard bootstrap argument that $k$ is of class $C^{\infty}$.
Next we show that $k$ satisfies (2.2) and (2.3). Let $\phi\in W^{2,2}(0,L)\cap W^{1,2}_{0}(0,L)$ be arbitrary and consider $\eta(x):=\phi(s(x))\mathbf{n}(s(x))$. With the help of integration by parts, we reduce (2.1) to
@@
Let $g:(\frac{1}{\sqrt{2}},1)\to\mathbf{R}$ be the function defined by (1.3). Then,
@@
for $q\in(\frac{1}{\sqrt{2}},1)$. In addition, the function $g$ has exactly two local extrema at the points $\hat{q}\simeq 0.79257$ and $q_{*}\simeq 0.90891$. More precisely, $g(\hat{q})=:\hat{\lambda}\simeq 0.70107$ is a unique local maximum and $g(q_{*})=0$ is a local minimum and $g$ is strictly monotone in $(\frac{1}{\sqrt{2}},\hat{q}]$, $[\hat{q},q_{*}]$ and $[q_{*},1)$.
@@
Let $\gamma\in A_{\ell}$ be a penalized pinned elastica and $k$ be the signed curvature of $\gamma$. Recall from Lemma 2.2 that $k$ satisfies (2.2) and (2.3). Then $k\equiv 0$ is a trivial solution to (2.2) and (2.3), so in the following we consider the other solutions. By [11, Proposition 3.3], any solution to the initial value problem for (2.2) with $k(0)=0$ is given by
@@
Note that $Y_{\sigma,\alpha,q}(L)=0$ follows from (2.15). Observe that by Lemma A.1 $02\mathrm{E}(q)-\mathrm{K}(q)>$ 0 if $q\in(0,q_{*})$ and $02\mathrm{E}(q)-\mathrm{K}(q)<$ 0 if $q\in(q_{*},1)$. Then, combining the boundary condition $\gamma(1)=\tilde{\gamma}(L)=(\ell,0)$ with (2.14) and (2.17), we obtain
@@
$\tilde{\gamma}(s)$ $=\begin{pmatrix}X_{\sigma,\alpha,q}(s)\\ Y_{\sigma,\alpha,q}(s)\end{pmatrix}\text{ with }q=q_{1}(\tfrac{\lambda\ell^{2}}{n^{2}})\ \text{ or }\ q_{2}(\tfrac{\lambda\ell^{2}}{n^{2}}),\quad$
@@
The curves $\gamma_{\rm sarc}^{\lambda,\ell,n}$, $\gamma_{\rm larc}^{\lambda,\ell,n}$, and $\gamma_{\rm loop}^{\lambda,\ell,n}$ obtained in Theorem 1.1 are $\frac{n}{2}$-fold well-periodic curves in terms of [19, Definition 2.6]. This fact will be used when we discuss the instability (see Subsection 3.2).
@@ 2.1. Properties of penalized pinned elasticae
$X(L-s)+X(s)=-\frac{4}{\alpha}\mathrm{E}(q)+\frac{2}{\alpha}\mathrm{K}(q)=\ell,\quad\textrm{for all }\ s\in[0,L].$
@@ 3.1. Stability of one-mode arcs
    $\Set{\gamma\in W^{2,2}_{\rm imm}(0,1;\mathbf{R}^{2})}{\gamma(0)=(0,0),\ \gamma(1)=(\ell,0),\ L[\gamma]=L[\gamma_{q}]}.$
@@
Using the above $\delta>0$, we now fix an arbitrary $\Gamma\in A_{\ell}$ with $\|\Gamma-\gamma_{q_{0}}\|_{W^{2,2}}<\delta$. This implies in particular that $|L[\Gamma]-L[\gamma_{q_{0}}]|<\delta$. Property (i) also implies that there exists $q\in(0,1)$ such that $L[\Gamma]=L[\gamma_{q}]$. From (3.3) follows that $q\in(q_{0}-\varepsilon,q_{0}+\varepsilon)$. Moreover, in view of property (ii), we see that $\mathcal{E}_{\lambda}[\Gamma]\geq\mathcal{E}_{\lambda}[\gamma_{q}]$, and this together with (3.2) yields that
@@
    $A_{\ell,q}:=\Set{\gamma\in W^{2,2}_{\rm imm}(0,1;\mathbf{R}^{2})}{\gamma(0)=(0,0),\ \gamma(1)=(\ell,0),\ L[\gamma]=L[\gamma_{w}(\cdot,q)]}.$
@@
Property (i) follows from the fact that $Q(q)=2\frac{\mathrm{E}(q)}{\mathrm{K}(q)}-1$ is continuous and strictly decreasing (cf. [20, Lemma B.4]), and satisfies $Q(0)=1$ and $Q(q_{*})=0$.
Next we show property (ii). The curve $\gamma_{w}(\cdot,q)$ coincides with $\hat{\gamma}^{-}_{0}$ in terms of [25, Theorem 1.1] since the modulus $q\in(0,q_{*})$ is uniquely determined by $\frac{\ell}{L}=2\frac{\mathrm{E}(q)}{\mathrm{K}(q)}-1$ and since the signed curvature of $\gamma_{w}(\cdot,q)$ is given by
@@
(see [25, proof of Theorem 1.1] for the coincidence of the signed curvature). Then it follows from [25, Theorem 1.3] that $\gamma_{w}(\cdot,q)=\hat{\gamma}^{-}_{0}$ is a minimizer of $B$ in $A_{\ell,q}$. ∎
@@ 3.2. Instability of higher modes ($n\geq 2$)
We apply the general rigidity principles obtained in [19] to deduce the instability of $\gamma_{\rm sarc}^{\lambda,\ell,n}$, $\gamma_{\rm larc}^{\lambda,\ell,n}$, and $\gamma_{\rm loop}^{\lambda,\ell,n}$ for $n\geq 2$. By Lemma 2.2 the bending energy $B$ satisfies [19, Hypotheses (H1’) and (H2)] with the choice of $\mathcal{F}=B$ and clearly satisfies [19, Hypothesis (H3)]. This fact together with Remark 2.8 allows us to apply [19, Theorems 2.3, 2.7, and 2.8] to the case of $\mathcal{F}=B$ and $\gamma=\gamma_{\rm sarc}^{\lambda,\ell,n}$, $\gamma_{\rm larc}^{\lambda,\ell,n}$, or $\gamma_{\rm loop}^{\lambda,\ell,n}$.
@@
Let $\gamma$ be either $\gamma_{\rm sarc}^{\lambda,\ell,n}$, $\gamma_{\rm larc}^{\lambda,\ell,n}$, or $\gamma_{\rm loop}^{\lambda,\ell,n}$ for some $n\geq 3$ and $L=L[\gamma]$. It follows from the formula of the signed curvature obtained in Theorem 1.1 that $k(0)=k(\frac{L}{n})=k(\frac{2L}{n})=0$, so that $\gamma$ satisfies [19, Assumption (2.2)]. Therefore, by [19, Theorem 2.3 and Remark 4.3] $\gamma$ is not a local minimizer of $B$ in
$A_{\ell,L}=\Set{\gamma\in W^{2,2}_{\rm imm}(0,1;\mathbf{R}^{2})}{\gamma(0)=(0,0),\ \gamma(1)=(\ell,0),\ L[\gamma]=L}.$
@@
(3.12) If γ is not a local minimizer of B in A,L,then γ is also not a local minimizer of λ in $\begin{split}&\text{If $\gamma$ is not a local minimizer of $B$ in $A_{\ell,L}$},\\ &\text{then $\gamma$ is also not a local minimizer of $\mathcal{E}_{\lambda}$ in $A_{\ell}$.}\end{split}$
@@
Since $\gamma_{\rm sarc}^{\lambda,\ell,2}$, $\gamma_{\rm larc}^{\lambda,\ell,2}$, and $\gamma_{\rm loop}^{\lambda,\ell,2}$ are $1$-fold well-periodic curves in the sense of [19, Definition 2.6] (recall Remark 2.8), we deduce from [19, Theorem 2.7] that $\gamma$ is not a local minimizer of $B$ in $A_{\ell,L}$. This fact together with (3.12) yields the desired conclusion. ∎

### v2

@@ 1. Introduction
where $k$ and $s$ respectively denote the signed curvature and the arclength parameter of $\gamma$. A critical point of the bending energy under the length-constraint is called Euler’s elastica, and it is known as a model of an elastic rod (cf. [APbook, LLbook, Love]).
@@
where $\lambda>0$ is called a penalization parameter, and $L[\gamma]=\int_{\gamma}\mathrm{d}s$, which is an object of interest in the recent literature (see e.g. [AGP20, Born, LS_JDG, Maddocks1981, Miura20, MW_24X, Sachkov08, Schrader16, Ydcds] for studies of critical points, [DP2014, Dia24, DKS2002, KM24, MPP21, OkabeNovaga2017] for gradient flows, and references therein).
@@
[…]
@@
Inspired by this question, we will obtain various properties of penalized pinned elasticae, such as (i) complete classification, (ii) stability results, (iii) energy comparison, and (iv) consequences for the elastic flow.
@@
Suppose that $\gamma\in A_{\ell}$ is a critical point of $\mathcal{E}_{\lambda}$ in $A_{\ell}$. Then, either $\gamma$ is a trivial line segment or $\gamma$ is, up to reflection and reparametrization, represented by one of the following
@@
Explicit formulae for each penalized pinned elasticae are also obtained (see Definition 2.7 below). Note that the classification strongly depends on the relation of $\lambda\ell^{2}$ and $\hat{\lambda}$ defined in (1.4). More precisely, if $\lambda\ell^{2}<\hat{\lambda}$, then $n_{\lambda,\ell}=1$ holds so that $\gamma_{\rm sarc}^{\lambda,\ell,1}$ and $\gamma_{\rm larc}^{\lambda,\ell,1}$ appear. On the other hand, if $\lambda\ell^{2}=\hat{\lambda}$, then $\gamma_{\rm sarc}^{\lambda,\ell,1}$ and $\gamma_{\rm larc}^{\lambda,\ell,1}$ coincide, and if $\lambda\ell^{2}>\hat{\lambda}$, then $n_{\lambda,\ell}\geq 2$ holds so that $\gamma_{\rm sarc}^{\lambda,\ell,1}$ and $\gamma_{\rm larc}^{\lambda,\ell,1}$ do not appear (see Figure 2). Such ‘bifurcation phenomena’ for elasticae have previously been found in e.g. [Lin98TAMS, Section 2.3.3-2.3.5] and [MW_24X, Remark 4.3].
@@
We also seek to compare the energies of the above critical points and study which one is energy-minimal (if one excludes the trivial line segment). As the explicit formulae show, $\mathcal{E}_{\lambda}[\gamma_{\rm larc}^{\lambda,\ell,1}]$ and $\mathcal{E}_{\lambda}[\gamma_{\rm loop}^{\lambda,\ell,1}]$ converge to 0 as $\lambda\to 0$. This implies that for small $\lambda>0$ it is not easy to determine which has less energy, $\gamma_{\rm larc}^{\lambda,\ell,1}$ or $\gamma_{\rm loop}^{\lambda,\ell,1}$. On the other hand, for larger $\lambda>0$ a comparison with $\gamma_{\rm sarc}^{\lambda,\ell,1}$ is also required. Indeed, since
@@
Here we are interested in the question of eventual embeddedness. We find a sharp energy threshold below which the above flow destroys any self-intersection in finite time. For more results on (not necessarily eventual) embeddedness of elastic flows we refer to [Bla10, MMR23].
@@
The value $C_{\lambda,\ell}$ is optimal in the sense that the above conclusion fails if we put any larger constant than $C_{\lambda,\ell}$ in (1.9). Indeed, there exists a smooth curve $\gamma_{0}$ with $\mathcal{E}_{\lambda}[\gamma_{0}]=C_{\lambda,\ell}$ and such that the $\lambda$-elastic flow with initial datum $\gamma_{0}$ is not embedded for all $t\geq 0$. Actually, one may choose $\gamma_{0}=\gamma_{\rm loop}^{\lambda,\ell,1}$.
The instability of $\gamma_{\rm loop}^{\lambda,\ell,1}$ in Theorem 1.2 allows us to construct a nonembedded smooth curve $\gamma_{0}\in A_{\ell}$ satisfying the condition (1.9), cf. Remark 5.3. From Theorem 1.4 we then infer that the flow destroys each self-intersection of $\gamma_{0}$ in finite time.
We close this introduction by comparing the properties of our penalized pinned elasticae with those already known from [MY_Crelle] about the critical points of $B$ in
@@
where $L>\ell$ is given. It turns out that the curves of Theorem 1.1 are also critical points in $A_{\ell,L}$ if $L$ is chosen to be their corresponding length. In [MY_Crelle], stability of these curves in $A_{\ell,L}$ is investigated. A major difference is that depending on the choice of $L$ it is possible that $\gamma_{\rm sarc}^{\lambda,\ell,1}$ is stable in $A_{\ell,L}$, whereas this is never the case for our penalized problem (cf. Theorem 1.2). Another difference is that in $A_{\ell,L}$ stable critical points are always unique whereas for our penalized problem the number of stable critical points depends on $\lambda$. Exposing these differences to the fixed-length problem is a main novelty of the present article.
@@ Acknowledgments
This work was initiated during the second author’s visit at Freiburg University. The second author is very thankful to the first author for his warm hospitality and providing a fantastic research atmosphere. Both authors are grateful to Tatsuya Miura for helpful suggestions. The authors are grateful to the referee whose suggestions have led to substantial improvements in this article. The second author is supported by FMfI Excellent Poster Award 2022 and JSPS KAKENHI Grant Number 24K16951.
@@ 2. Classification
    For every smooth family $\{\gamma_{\varepsilon}\}_{\varepsilon\in(-\varepsilon_{0},\varepsilon_{0})}\subset A_{\ell}$ with $\varepsilon_{0}>0$ such that $\gamma_{0}=\gamma$, it follows that
    (2.1) $\frac{d}{d\varepsilon}\mathcal{E}_{\lambda}[\gamma_{\varepsilon}]\Big|_{\varepsilon=0}=0.$
@@
First we show that $k$ is smooth on $(0,L)$. For arbitrary $\varphi\in W^{2,2}(0,L;\mathbf{R}^{2})\cap W^{1,2}_{0}(0,L;\mathbf{R}^{2})$ we consider $\eta(x):=\varphi(s(x))$ for $x\in[0,1]$, where $s$ denotes the arclength function, i.e., $s(x):=\int_{0}^{x}|\gamma^{\prime}|$. Choosing $\gamma_{\varepsilon}=\gamma+\varepsilon\eta$ in (2.1), and using the known formulae of first derivative of $B$ and $L$ (cf. [MY_AMPA, Lemma A.1]), we deduce that
@@
where $\mathbf{t}$ and $\mathbf{n}$ denote the unit tangent vector and the unit normal vector of $\tilde{\gamma}$, respectively. Since this identity holds in particular for all $\varphi\in C^{\infty}_{\rm c}(0,L;\mathbf{R}^{2})$, we can deduce from a standard bootstrap argument that $k$ is of class $C^{\infty}(0,L)$. More careful analysis of the equation, which is by now standard, shows also that $k\in C^{\infty}([0,L])$.
Next we show that $k$ satisfies (2.2) and (2.3). Let $\phi\in W^{2,2}(0,L)\cap W^{1,2}_{0}(0,L)$ be arbitrary and consider $\eta(x):=\phi(s(x))\mathbf{n}(s(x))$ and again look at $\gamma_{\varepsilon}:=\gamma+\varepsilon\eta$. With the help of integration by parts, we reduce (2.1) to
@@
Let $g:[\frac{1}{\sqrt{2}},1)\to\mathbf{R}$ be the function defined by (1.3). Then,
@@
for $q\in(\frac{1}{\sqrt{2}},1)$. In addition, the function $g$ has exactly two local extrema at the points $\hat{q}\simeq 0.79257$ and $q_{*}\simeq 0.90891$. More precisely, $g(\hat{q})=:\hat{\lambda}\simeq 0.70107$ is a unique local maximum and $g(q_{*})=0$ is a local minimum and $g$ is strictly monotone in $(\frac{1}{\sqrt{2}},\hat{q}$], $\hat{q},q_{*}]$ and $q_{*},1)$.
@@
Let $\gamma\in A_{\ell}$ be a penalized pinned elastica and $k$ be the signed curvature of $\gamma$. Recall from Lemma 2.2 that $k$ satisfies (2.2) and (2.3). Then $k\equiv 0$ is a trivial solution to (2.2) and (2.3), so in the following we consider the other solutions. By [Lin96, Proposition 3.3], any solution to the initial value problem for (2.2) with $k(0)=0$ is given by
@@
Note that $Y_{\sigma,\alpha,q}(L)=0$ follows from (2.15). Observe that by Lemma A.1 $2\mathrm{E}(q)-\mathrm{K}(q)>0$ if $q\in(0,q_{*})$ and $2\mathrm{E}(q)-\mathrm{K}(q)<0$ if $q\in(q_{*},1)$. Then, combining the boundary condition $\gamma(1)=\tilde{\gamma}(L)=(\ell,0)$ with (2.14) and (2.17), we obtain
@@
$\tilde{\gamma}(s)$ $=\begin{pmatrix}X_{\sigma,\alpha,q}(s)\\ Y_{\sigma,\alpha,q}(s)\end{pmatrix}\text{ with }q=q_{1}(\tfrac{\lambda\ell^{2}}{n^{2}})\ \text{ or }\ q_{2}(\tfrac{\lambda\ell^{2}}{n^{2}}),\quad\text{or}\quad$
@@
The curves $\gamma_{\rm sarc}^{\lambda,\ell,n}$, $\gamma_{\rm larc}^{\lambda,\ell,n}$, and $\gamma_{\rm loop}^{\lambda,\ell,n}$ obtained in Theorem 1.1 are $\frac{n}{2}$-fold well-periodic curves in terms of [MY_Crelle, Definition 2.6]. This fact will be used when we discuss the instability (see Subsection 3.2).
@@ 2.1. Properties of penalized pinned elasticae
[…]
@@ 3.1. Stability of one-mode arcs
[…]
@@
Using the above $\delta>0$, we now fix an arbitrary $\Gamma\in A_{\ell}$ with $\|\Gamma-\gamma_{q_{0}}\|_{W^{2,2}}<\delta$. This implies in particular that $|L[\Gamma]-L[\gamma_{q_{0}}]|<\delta$. Property (i) also implies that there exists $q\in(0,1)$ such that $L[\Gamma]=L[\gamma_{q}]$. From (3.3) follows that $q\in(q_{0}-\epsilon,q_{0}+\epsilon)$ Moreover, in view of property (ii), we see that $\mathcal{E}_{\lambda}[\Gamma]\geq\mathcal{E}_{\lambda}[\gamma_{q}]$, and this together with (3.2) yields that
@@
[…]
@@
Property (i) follows from the fact that $Q(q)=2\frac{\mathrm{E}(q)}{\mathrm{K}(q)}-1$ is continuous and strictly decreasing (cf. [MR23, Lemma B.4]), and satisfies $Q(0)=1$ and $Q(q_{*})=0$.
Next we show property (ii). The curve $\gamma_{w}(\cdot,q)$ coincides with $\hat{\gamma}^{-}_{0}$ in terms of [Ydcds, Theorem 1.1] since the modulus $q\in(0,q_{*})$ is uniquely determined by $\frac{\ell}{L}=2\frac{\mathrm{E}(q)}{\mathrm{K}(q)}-1$ and since the signed curvature of $\gamma_{w}(\cdot,q)$ is given by
@@
(see [Ydcds, proof of Theorem 1.1] for the coincidence of the signed curvature). Then it follows from [Ydcds, Theorem 1.3] that $\gamma_{w}(\cdot,q)=\hat{\gamma}^{-}_{0}$ is a minimizer of $B$ in $A_{\ell,q}$. ∎
@@ 3.2. Instability of higher modes ($n\geq 2$)
We apply the general rigidity principles obtained in [MY_Crelle] to deduce the instability of $\gamma_{\rm sarc}^{\lambda,\ell,n}$, $\gamma_{\rm larc}^{\lambda,\ell,n}$, and $\gamma_{\rm loop}^{\lambda,\ell,n}$ for $n\geq 2$. By Lemma 2.2 the bending energy $B$ satisfies [MY_Crelle, Hypotheses (H1’) and (H2)] with the choice of $\mathcal{F}=B$ and clearly satisfies [MY_Crelle, Hypothesis (H3)]. This fact together with Remark 2.8 allows us to apply [MY_Crelle, Theorems 2.3, 2.7, and 2.8] to the case of $\mathcal{F}=B$ and $\gamma=\gamma_{\rm sarc}^{\lambda,\ell,n}$, $\gamma_{\rm larc}^{\lambda,\ell,n}$, or $\gamma_{\rm loop}^{\lambda,\ell,n}$.
@@
Let $\gamma$ be either $\gamma_{\rm sarc}^{\lambda,\ell,n}$, $\gamma_{\rm larc}^{\lambda,\ell,n}$, or $\gamma_{\rm loop}^{\lambda,\ell,n}$ for some $n\geq 3$ and $L=L[\gamma]$. It follows from the formula of the signed curvature obtained in Theorem 1.1 that $k(0)=k(\frac{L}{n})=k(\frac{2L}{n})=0$, so that $\gamma$ satisfies [MY_Crelle, Assumption (2.2)]. Therefore, by [MY_Crelle, Theorem 2.3 and Remark 4.3] $\gamma$ is not a local minimizer of $B$ in
[…]
@@
[…]
@@
Since $\gamma_{\rm sarc}^{\lambda,\ell,2}$, $\gamma_{\rm larc}^{\lambda,\ell,2}$, and $\gamma_{\rm loop}^{\lambda,\ell,2}$ are 11-fold well-periodic curves in the sense of [MY_Crelle, Definition 2.6] (recall Remark 2.8), we deduce from [MY_Crelle, Theorem 2.7] that $\gamma$ is not a local minimizer of $B$ in $A_{\ell,L}$. This fact together with (3.12) yields the desired conclusion. ∎
@@
Remark 3.9.
While $\gamma_{\rm loop}^{\lambda,\ell,1}$ is unstable as in Theorem 1.2, it will be shown in Theorem 1.3 that $\gamma_{\rm loop}^{\lambda,\ell,1}$ is the (unique) minimizer among penalized pinned elasticae except for a trivial global minimizer. Notice carefully that these results are not contradictory. In fact, minimizing a functional among a subset of its critical points is different from minimizing the functional in general. Since perturbations of critical points are not necessarily critical points, approaches like stability analyses can not be used to address the question of minimality among critical points.
Next we examine the energy landscape in the neighborhood of $\gamma_{\rm loop}^{\lambda,\ell,1}$. We can find not only an energy-decreasing perturbation (as in Lemma 3.5) but also and an energy-increasing perturbation as follows. Let $\gamma\in A_{\ell}$ be the reparametrization of $\gamma_{\rm loop}^{\lambda,\ell,1}$, and $0<a<b<1$ be such that $\gamma(a)=\gamma(b)$, i.e. the self-intersection point. For $|\varepsilon|<1$, define $\gamma_{\varepsilon}:[0,1]\to\mathbf{R}^{2}$ by the constant-speed reparametrization of
$\begin{cases}(\varepsilon+1)\big(\gamma(x)-\gamma(a)\big)+\gamma(a)\quad&x\in[a,b],\\ \gamma(x)&\text{otherwise},\end{cases}$
i.e., constructed by dilation of the loop. Setting
$G(\varepsilon):=\frac{1}{\varepsilon+1}B[\gamma|_{[a,b]}]+(\varepsilon+1)\lambda L[\gamma|_{[a,b]}],$
we can compute the energy gap as $\mathcal{E}_{\lambda}[\gamma_{\varepsilon}]-\mathcal{E}_{\lambda}[\gamma]=G(\varepsilon)-G(0)$. Thus the second variation along this perturbation is given by $G^{\prime\prime}(0)=2B[\gamma|_{[a,b]}]>0$. Notice also that the first variation vanishes as $\gamma_{\rm loop}^{\lambda,\ell,1}$ is a critical point. Hence, one can find that $\gamma$ is a local minimizer along this direction. This result together with the existence of a perturbation satisfying (3.8) imply that $\gamma_{\rm loop}^{\lambda,\ell,1}$ behaves like a saddle point.